OpenStudy (anonymous):

Decide how many solutions this equation has: x^2 + 3 = 0

OpenStudy (anonymous):

@twistnflip

ganeshie8 (ganeshie8):

may be just try solving it : \(x^2+3=0 \implies x^2=-3 \implies \cdots \)

OpenStudy (anonymous):

Okay can you explain to me how to solve it. I came uo with x=6 so one solution but i dont know if thats correct

ganeshie8 (ganeshie8):

plugin \(x=6\) does it make the given equation true ?

OpenStudy (anonymous):

I think not

ganeshie8 (ganeshie8):

x^2 + 3 = 0 6^2 + 3 = 0 36 + 3 = 0 39 = 0 which is clearly false, so nope, x=6 is not a solution.

OpenStudy (anonymous):

What about a square root of negative 3? Does such thing exist?

OpenStudy (anonymous):

I think there is no solution right?

OpenStudy (anonymous):

What if I try putting the negative sign before the bracket surrounding square root of 3?

OpenStudy (anonymous):

Like -(sqrt(3)) Probably I am wrong.

OpenStudy (anonymous):

Nah forget it I don't think that even exists.

OpenStudy (anonymous):

So yeah, I think x would have to be of square root value but with the negative sign I think that makes the equation unsolvable. Is that correct?

OpenStudy (freckles):

depends on the set we are solving over but no sqrt(-3) doesn't equal -sqrt(3) sqrt(-3) is imaginary while -sqrt(3) is real

OpenStudy (anonymous):

Right. Imaginary root shuold be used as the solution. So I belive minus square root of three might be more relevant? Forgive me if I am wrong.

OpenStudy (freckles):

\[(-\sqrt{3})^2=((-1)\sqrt{3})^2= (-1)^2 (\sqrt{3})^2=1(3)=3 \neq -3\]

OpenStudy (anonymous):

So that's the full solution for x apparently. @abriannaridley I guess that's your take home message from todayXD

OpenStudy (anonymous):

lol thanks for everyones help! :P

OpenStudy (freckles):

what conclusion did you draw from this @abriannaridley ?

ganeshie8 (ganeshie8):

im liking jokes that cortana tells me when i feel bored... so let me tell you a joke :) gimme one moment to type..

OpenStudy (anonymous):

okay so im not 100% if this is right but there is only solution to that problem? @freckles

OpenStudy (freckles):

only solution you mean no real solution right?

OpenStudy (freckles):

x^2=negative number will never happen under the real numbers

OpenStudy (anonymous):

Does cortana make no comment like Siri when you talk like a you know person ?

OpenStudy (anonymous):

Yes lol! should've just said no solution

OpenStudy (freckles):

x^2=negative number can happen under the complex numbers

OpenStudy (anonymous):

Find the x-intercepts of the graph of the equation: y = x2 - 3x + 2 im not sure how to do this one either should i just post it as a new question?

OpenStudy (freckles):

x-intercepts can found by setting y to 0 and solve for x

OpenStudy (anonymous):

ohh okay!

OpenStudy (freckles):

and yeah to prevent things from getting to long sometimes it is best to post a new question

OpenStudy (freckles):

but you can post a question I guess if you have trouble solving 0=x^2-3x+2

ganeshie8 (ganeshie8):

sry for the delay, here it is : ``` A foreman is yelling down into an excavation : "How many of you men down there ?" the reply came back, "three." and he said "Okay, half of you come on up." ```

OpenStudy (freckles):

but you can post a question I guess if you have trouble solving 0=x^2-3x+2

ganeshie8 (ganeshie8):

that joke refers to solving the equation \(2x=3\). does it really have a solution ? If it has a solution, what does it represent ?